Hemoglobin is the oxygen transport protein in the red blood cells of all vertebrates. It's emperical formula is C738H1166N203O208S2Fe. It's molecular mass is 65332.91 amu, what is the molecular formula of hemoglobin?

How do I solve this problem?
Answer #1

First,

calculate the empirical formula mass

C738H1166N203O208S2Fe

C = 738

H= 1166

N = 203

O = 208

S = 2

Fe = 1

then, calculate all masses; note that MW states for molar weight

MW of C = 12; MW of H = 1; MW of N = 14; MW of O = 16; MW of S = 32; MWof Fe = 55.8

then

total mass = addition of all masses

total mass = (12)*(738) + 1*(1166) + 14*203+ 16*208 + 32*2 + 55.8*1 = 16311.8 g/mol

since we know that the total MW of hemoglobin is = 65332.91 g/mol

then

find how many times does the empirical formula "fits" in the total MW

n = MW total / MW empirical = 65332.91 /16311.8 = 4.005254

n = 4

then

C738H1166N203O208S2Fe --> 4

then

C738*4 H1166*4 N203*4 O208*4 S2*4 Fe*4

C2952H4664N812O832S8Fe4